OpenStudy (anonymous):

Free medal and fan for answer :) What is the slope of the line that passes through the points (-3, 15) and (-9, 7)?

OpenStudy (anonymous):

slope of the line passes through the points (x1,y1) and (x2,y2) is \[\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (anonymous):

I don't get it

OpenStudy (yanasidlinskiy):

\[\frac{ 7+3 }{ -9-15 }\]

OpenStudy (yanasidlinskiy):

Got it?

OpenStudy (yanasidlinskiy):

\[\frac{ 10 }{ 24 }\]

OpenStudy (yanasidlinskiy):

Got to keep going.

OpenStudy (anonymous):

4/3?

OpenStudy (yanasidlinskiy):

Did you continue from the fraction of what I gave you?

OpenStudy (yanasidlinskiy):

@Emllyk ?

OpenStudy (anonymous):

i think so...

OpenStudy (yanasidlinskiy):

Look here: \[\frac{ 10 }{ 24 } \div \frac{ 2 }{ 2 } = \frac{ 5 }{ 12 }\]